Skip to Search

 
#

Support Services Meeting Notes
March 6, 2006

Present: Bill Covey, Brian Keith, Rob Roberts, Grace Strawn
Systems
  1. NT2R - is down. He is trying to get it back up under another name. The cluster node that manages the name NT2R thinks there is a duplicate on the network. Will changed the name and it came back online, then it later failed. Will is working on it.
  2. Facilities is reporting Outlook outages. Could be a Cox cable problem? Need to determine if the Circulation folks are having difficulty. Cox did mention they were scheduled to do an IPOP change. May not have fixed the routing table.
  3. Bobbie out for the next month. Geoffrey is out today.
Library HR
  1. Brian has conducted five exit interviews in the past week.
  2. Performance Training sessions - Trudi is tracking attendance. Evaluations are due back March 27. If anyone is to fall in the last two categories, UF HR will have to get involved. Has to be legit and arguable to UF. It is a formal, lengthy process. Documentation on all employees to be kept by Library HR.
Facilities
  1. Systems room 19D (server/tape room) - lighting is complete.
  2. VOIP - Education will be done in the next couple of weeks.
  3. Rob met with Carol Drum and they put together a list as to what work needs to be done in MSL. Rob suggests that rather than painting, using a commercial grade wallpaper.
  4. Friday we got an estimate for carpeting the Systems area. Bill says it won't happen this fiscal year. Have yet to figure out the status of carry-forward funds.
  5. Supply estimate was given to Business Services by Russ. Grace had some questions. Ongoing.
  6. Shed roof not finished yet.
  7. 11x14 scanner lost - DLC. No one knows where it is. All paperwork indicates it should be in DLC. Systems will re-check to make sure it was not replaced. Otherwise, Facilities is planning to file a police report.
Business Services
  1. Supply requests - amount to get us through until the next fiscal year - not as much as expected.
  2. Dale - MG&G - no money. Incorrect info - money added to wrong account. Have to transfer funds.
  3. OPS checks - emergency check had to be cut for Documents student. Supervisor was unable to approve the students' time. Situation now rectified.
  4. Barbara Oliver is out.
  5. Job description for Dolapo to be sent to Library HR today.
Directors
  1. Budget panic - due today to the Provost (written version). The presentation will be 2-3 days hence. We'll have a big request - asking them to cover the serials cut. The Provost is at least aware of the cut, as it mentioned in a resent faculty meeting. It will be noted that we are $3.7 million behind other top 10 libraries. We are going to make our points as best we can and see how they respond. Money for positions - need five new slots in order to stay open 24-5. Two more needed to preserve retrieval service. We are looking at expanding this popular service.
  2. West - we've got three installers going right now. Rob mentions that he hopes the sixth floor will be open this week as Facilities has been storing a lot of equipment (chairs, tables, gates) and would like to get it into West. Per Bill, the sixth floor is not an option right now - maybe the fourth. The real problem is the insurance issue.
  3. Butler Properties - demo to sell property - we have some legal issues to tie up with them. Need to certify that we have not subleased and they don't owe us.
  4. MSL compact shelving - trying to get it back up - Space Saver is going to come by and take a look at it. Right now they are having to crank it by hand.

 

University of Florida Home Page